Shift report 20011025 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Oct 25 2001 - 23:53:30 EDT

  • Next message: Ian Bearden: "RE: TOFW Pid"

    ------------------------------------------------------------
    Date: Thu Oct 25 23:53:30 2001, Shift: 16:00-24:00
    Supervisor: DO, with: JC, CEJ
    Log book#10 pages 56 - 57
    
    Shift Summary:
    ------------------------------------------------------------
    1.Machine status: bad
      QLI during most of the shift  
      MCR made some beam tests. We had beam for 20 mn
      but the single ZDC rates were really high compared
      to coincidences. It was not possible to take data 
      under these conditions. MCR dumped the beams and since
      then nothing.
    
    2.No achievements at all.
    
    3.DAQ problem around 17:00, could not talk to FB3, 
      therefore could not take data. Flemming had to reboot
      it (which doesn't happen quite often) since the usual
      tricks did not work. Made a small test, successful. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Oct 25 2001 - 23:54:02 EDT